solve the equation:
1. 4 sin^-1 (x) = π
2. 3 cos^-1 (2x) = 2π
3. 3 tah^-1 x = π
4. 4 cos^-1 x - 2π = 2 cos^-1 x

Answer by lwsshak3(11628)

solve the equation:
1. 4 sin^-1 (x) = π
sin^-1 (x) = π/4
x=sin(π/4)=√2/2
..
2. 3 cos^-1 (2x) = 2π
cos^-1 (2x) = 2π/3
2x=cos(2π/3)
x=cos(2π/3)/2=(-1/2)/2=-1/4
..
3. 3 tan^-1 x = π
tan^-1(x)=π/3
x=tan(π/3)=√3
..
4. 4 cos^-1 x - 2π = 2 cos^-1 x
2 cos^-1 (x)=2π
cos^-1 (x)=π
x=cosπ=-1